# Review of Generative AI Methods in Cybersecurity

### 13 Mar 2024 v1

@@ 4 Cyber Offense
Gupta et al. [22] have shown that ChatGPT could create social engineering attacks, phishing attacks, automated hacking, attack payload generation, malware creation, and polymorphic malware. Experts might be motivated to automate numerous frameworks, standards, and guidelines (Figure 8) to use GenAI for security operations. However, the end products can also be utilised for offensive cyber operations. This not only increases the pace of attacks but also makes attribution harder. An attribution project typically utilizes frameworks like the MICTIC framework, which involves the analysis of Malware, Infrastructure, Command and Control, Telemetry, Intelligence, and Cui Bono[28]. Many behavioural patterns for attribution, such as code similarity, compilation timestamps, working weeks, holidays, and language, could disappear when Gen AI creates Offensive Cyber Operations (OCO) code. This makes attribution more challenging, especially if the whole process becomes automated.
@@ 4.1 Social engineering
Falade [29] investigates the application of generative AI in social engineering, assuming the definition of social engineering as an array of tactics employed by adversaries to manipulate individuals into divulging confidential information or performing actions that may compromise security. The study underscores tools like ChatGPT, FraudGPT, and WormGPT in enhancing the authenticity and specificity of phishing expeditions, pretexting, and the creation of deepfakes. The author reflects on the double-edged impact of advancements like Microsoft’s VALL-E and image synthesis models like DALL·E 2, drawing a trajectory of the evolving threat landscape in social engineering through deepfakes and exploiting human cognitive biases.
@@ 4.3 Automated hacking
PentestGPT[31], or GPTs[32](custom versions of ChatGPT that can be created for a specific purpose) like GP(en)T(ester)[33]. Pentest Reporter[34] are introduced as applications built on ChatGPT, designed to assist in penetration testing—a sanctioned simulation of cyberattacks on systems to evaluate security. However, these tools could also be adapted for malicious purposes in automated hacking. Many emerging tools, such as WolfGPT, XXXGPT, and WormGPT, have been invented; however, no study has yet evaluated and compared their real offensive capabilities. Gupta et al.[22] noted that an AI model could scan new code for similar weaknesses with a comprehensive dataset of known software vulnerabilities, pinpointing potential attack vectors. While AI-assisted tools like PentestGPT are intended for legitimate and constructive uses, there is potential for misuse by malicious actors who could create similar models to automate unethical hacking procedures. If fine-tuned to identify vulnerabilities, craft exploitation strategies, and execute those strategies, these models could potentially pose significant threats to cybersecurity. However, this enormous task should be divided into smaller segments, such as reconnaissance, privilege escalation, etc. Temara[35] outlines how ChatGPT can be utilized during the reconnaissance phase by employing a case study methodology to demonstrate collecting reconnaissance data such as IP addresses, domain names, network topologies, and other critical information like SSL/TLS cyphers, ports and services, and operating systems used by the target. Happe et al. [36] investigate the use of Large Language Models (LLMs) in Linux privilege escalation. The authors introduce a benchmark for automated testing of LLMs’ abilities to perform privilege escalation using a variety of prompts and strategies. They implement a tool named Wintermute, a Python program that supervises and controls the privilege-escalation attempts to evaluate different models and prompt strategies. Their findings indicate that GPT-4 generates the highest quality commands and responses. In contrast, Llama2-based models struggle with command parameters and system descriptions. In some scenarios, GPT-4 achieved a 100% success rate in exploitation.
@@ 4.4 Attack payload generation
Studies [22, 37] have highlighted the capacity of Large Language Models (LLMs), particularly ChatGPT, for payload generation. Our examination of GPT-4’s current abilities confirmed its proficiency in generating payloads and embedding them into PDFs (as an example) using a reverse proxy(Figure 9). The following is a summation of the frameworks GPT-4 utilizes with successful payload code generation, accompanied by their respective primary functions:
@@ 4.5 Malware code generation
Gupta et al.[22] mentioned they could obtain potential ransomware code examples by utilizing a ’DAN’ jailbreak. We tested all the existing DAN techniques outlined in [38]. At the time of our research, these techniques were no longer functional; therefore, we could not reproduce samples of WannaCry, Ryuk, REvil, or Locky, as addressed by [22]. However, we generated an educational ransomware code (Figure 10), applying basic code obfuscation(Renaming and Control Flow Flattening). ChatGPT has garnered significant attention from the cybersecurity community, leading to the implementation of robust filters. Nonetheless, this does not imply that other models, such as the Chinese 01.ai[39], will have an equivalent opportunity to mitigate the potential for misuse in generating malicious code.
@@ 4.6 Polymorphic malware
The usage of LLMs could see the rise of malware, which integrates improved evasion techniques and polymorphic capabilities [40]. This often relates to overcoming both signature detection and behavioural analysis. An LLM-based malware agent could thus focus on rewriting malware code which could change the encryption mode used or produce obfuscated code which is randomized for each build [41]. Gupta et al. [22] outlined a method of getting ChatGPT to seek out target files for encryption and thus mimic ransomware behaviour, but where it mutated the code to avoid detection. They even managed to embed a Python interpreter in the malware and where it could query ChatGPT for new software modules.
@@ 5 Cyber Defence
In the ever-evolving cybersecurity battlefield, the “Cyber Defence” segment highlights the indispensable role of GenAI in fortifying digital fortresses against increasingly sophisticated cyber threats. This section is dedicated to exploring how GenAI technologies, renowned for their advanced capabilities in data analysis and pattern recognition, are revolutionizing the approaches to cyber defence. Iturbe et al. [44] outline the AI4CYBER framework, which provides a roadmap for the integration of AI into cybersecurity applications. This includes AI4VUN (AI-enhanced vulnerability identification); AI4FIX (AI-driven self-testing and automatic error correction); AI4SIM (Simulation of advanced and AI-powered attacks); AI4CTI A(I-enhanced cyber threat intelligence of adversarial AI); AI4FIDS (federated learning-enhanced detection); AI4TRIAGE (Root cause analysis and alert triage); AI4SOAR (Automatic orchestration and adaptation of combined responses); AI4ADAPT (Autonomy and optimization of response adaptation); and AI4DECEIVE (Smart deception and honeynets); and AI4COLLAB (Information sharing with privacy and confidentiality).
@@ 5.2 Cybersecurity Reporting
Using LLMs provides a method of producing Cyber Threat Intelligence (CTI) using Natural Language Processing techniques. For this, Perrina et al. [47] created the AGIR (Automatic Generation of Intelligence Reports) system and which aims to link together text data from many data sources. For this, they found that AGIR has a high recall value (0.99) without any hallucinations, along with a high score of the Syntactic Log-Odds Ratio (SLOR).
@@ 5.3 Threat Intelligence
Bayer et al. [48] address the challenge of information overload in the gathering of Cyber Threat Intelligence (CTI) from open-source intelligence (OSINT). A novel system is introduced, utilizing transfer learning, data augmentation, and few-shot learning to train specialized classifiers for emerging cybersecurity incidents. In parallel, Microsoft Security Copilot [49] has been providing CTI to its customers using GPT, and operational use cases have been observed, such as the Cyber Dome initiative in Israel [50].
@@ 5.4 Secure Code Generation and Detection
Machine learning in code analysis for cybersecurity has been elaborated very well [51]. Recent progress in natural language processing (NLP) has given rise to potent language models like the GPT (Generative Pre-trained Transformer) series, encompassing large language models (LLM) like ChatGPT and GPT-4 [52]. Traditionally, Static Application Security Testing (SAST) is a method that employs Static Code Analysis (SCA) to detect possible security vulnerabilities. We are interested in seeing whether SAST or GPT could be more efficient in decreasing the window of vulnerability. The window of vulnerability is defined as when the most vulnerable systems apply the patch minus the time an exploit becomes active. The precondition is met if two milestones that assume the detection of vulnerabilities verify their effectiveness, along with the vendor patch [53].
Laws in some countries, like China, ban the reporting on zero-days (see articles 4 and 9 of [54]), and contests like the Tianfu Cup [55], which is a systematic effort to find zero days, proliferate zero-day discovery continuously. Therefore, this precondition may not be satisfied timely, especially if the confirmation of vulnerabilities is not verified. A wide window of vulnerability threatens national security if a zero-day has been taken against critical infrastructures. DARPA introduces an important challenge that may help overcome this threat: (AIxCC) [56]). Moreover, this topic touches a part of the BSI studies [57, 58], and where we can define two main classifications of software testing for cybersecurity bugs as:
  • Static Application Security Testing (SAST). This is often called White Box Testing, is a set of algorithms and techniques used for analyzing source code. It operates automatically in a non-runtime environment to detect vulnerabilities such as hidden errors or poor source code during development.
  • Dynamic Application Security Testing (DAST). This follows the opposite approach and analyzes the program while it is operating. Functions are called with values in the variables as each line of code is checked, and possible branching scenarios are guessed. Currently, GPT-4 and other LLMs can’t provide DAST capabilities because the code needs to run within the runtime for this to work, requiring many deployment considerations.
@@ 5.5 Vulnerability detection and repair
Codex, introduced by Mark et al. [63], represents a significant advancement in GPT language models, tailored specifically for code synthesis using data from GitHub. This refined model underpins the operations of GitHub Copilot. When assessed on the HumanEval dataset, designed to gauge the functional accuracy of generating programs based on docstrings, Codex achieved a remarkable f28.8% success rate. In stark contrast, GPT-3 yielded a 0% success rate, and GPT-J achieved 11.4%. A standout discovery was the model’s enhanced performance through repeated sampling, with a success rate soaring to 70.2% when given 100 samples per problem. Despite these promising results, Codex does exhibit certain limitations, notably struggling with intricate docstrings and variable binding operations. The paper deliberates on the broader ramifications of deploying such potent code-generation tools, touching upon safety, security, and economic implications.
In a technical evaluation, Cheshkov et al. [64] found that the ChatGPT and GPT-3 models, despite their success in various other code-based tasks, performed on par with a dummy classifier for this particular challenge. Utilizing a dataset of Java files sourced from GitHub repositories, the study emphasized the models’ current limitations in the domain of vulnerability detection. However, the authors remain optimistic about the potential of future advancements, suggesting that models like GPT-4, with targeted research, could eventually make significant contributions to the field of vulnerability detection.
@@
5.6 Evaluating LLMs for code security
The OWASP top 10 for LLMs [66] introduced ten security risks as follows: Prompt Injection, Insecure Output Handling, Training Data Poisoning, Model Denial of Service, Supply Chain Vulnerabilities, Sensitive Information Disclosure, Insecure Plugin Design, Excessive Agency, Overreliance, and Model Theft.
Elgedawy et al. [67] analysed the ability of LLM to produce both secure and insecure code and conducted experiments using GPT-3.5, GPT-4, Google Bard and Google Gemini from Google. This involved nine basic tasks in generating code and assessing for functionality, security, performance, complexity, and reliability. They found that Bard was less likely to link to external libraries, and thus be less exposed to software chain issues. There were also variable levels of security and code integrity, such as input validation, sanitization, and secret key management, and while useful for automated code reviews, LLMs often require manual reviews, especially in understanding the context of the deployed code. For security, GPT-3.5 seemed to be more robust for error handling and secure coding practices when there is security consciousness applied to the prompt, there was a lesser focus on this with GPT-4, but where there were more advisory notes given. Overall, Gemini produced the most code vulnerabilities, and the paper advised users to be careful when deploying secure code from Gemini.
@@ 5.7 Developing Ethical guidelines
Kumar et al. [68] outlined the ethical challenges related to LLMs and where the datasets that they were trained on could be open to breaches of confidentiality, including five major threats: prompt injection, jailbreaking, Personal Identifiable Information (PII) exposure, sexually explicit content, and hate-based content. They propose a model that provides an ethical framework for scrutinizing the ethical dimensions of an LLM within the testing phase. The MetaAID framework [69] focuses on strengthening cybersecurity using Metaverse cybersecurity Q&A and attack simulation scenarios, along with addressing concerns around the ethical implications of user input. The framework is defined across five dimensions of:
  • Ethics. This defines an alignment with accepted moral and ethical principles.
  • Legal Compliance. This defines that any user input does not violate laws and/or regulations. This might relate to privacy laws and copyright protection.
  • Transparency. This defines that user inputs must be clear in requirements, and does not intend to mislead the LLM.
  • Intent Analysis. This defines that user input should not have other intents, such as jailbreaking the LLM.
  • Malicious intentions. This defines that user input should be free of malicious intent, such as to perform hate crime.
  • Social Impact. This defines how user input could have a negative effect on society, such as searching for ways to do harm to others, such as related to crashing the stock market or planning a terrorist attack.
@@ 5.8 Incident Response and Digital Forensics
Scanlon et al. [16] investigated using a pre-trained LLM for artefact understanding, evidence searching, code generation, anomaly detection, incident response, and education. For this, the low-risk applications and many others still require expert knowledge. The key areas of strength include creativity, reassurance, and avoidance of the blank page syndrome.
Scanlon et al. [16] investigated using a pre-trained LLM for artefact understanding, evidence searching, code generation, anomaly detection, incident response, and education. For this, the low-risk applications and many others still require expert knowledge. The key areas of strength include creativity, reassurance, and avoidance of the blank page syndrome, especially in places where ChatGPT cannot get wrong, such as in forensic scenario creation and reassurance of evidence. Still, care must be taken in this to avoid ChatGPT hallucinations. Another practical application is in code generation and explanations, such as generating commands for tool integration - which can be used as a starting point in an investigation.
For weaknesses, Scanlon found that it was essential to have a good quality and up-to-date training model; otherwise, ChatGPT could be biased and outdated in its analysis. Generally, it might not be able to find the newest artefacts - if it is trained on relatively old data. Additionally, ChatGPT’s accuracy reduces as the task becomes more specific, and any analysis of non-textural data - such as network packets - is less accurate. The length of some evidence logs, too, caused problems and often had to be prefiltered before they were analysed. A final problem identified is that the output of ChatGPT is often not deterministic - which is unsuitable for reproducibility.

### 19 Mar 2024 v2

@@ 3 Cyber Offense
Gupta et al. [32] have shown that ChatGPT could create social engineering attacks, phishing attacks, automated hacking, attack payload generation, malware creation, and polymorphic malware. Experts might be motivated to automate numerous frameworks, standards, and guidelines (Fig. 9) to use GenAI for security operations. However, the end products can also be utilised for offensive cyber operations. This not only increases the pace of attacks but also makes attribution harder. An attribution project typically utilizes frameworks like the MICTIC framework, which involves the analysis of Malware, Infrastructure, Command and Control, Telemetry, Intelligence, and Cui Bono [38]. Many behavioural patterns for attribution, such as code similarity, compilation timestamps, working weeks, holidays, and language, could disappear when GenAI creates Offensive Cyber Operations (OCO) code. This makes attribution more challenging, especially if the whole process becomes automated.
@@ 3.1 Social engineering
Falade [39] investigates the application of generative AI in social engineering, assuming the definition of social engineering as an array of tactics employed by adversaries to manipulate individuals into divulging confidential information or performing actions that may compromise security. The study underscores tools like ChatGPT, FraudGPT, and WormGPT to enhance the authenticity and specificity of phishing expeditions, pretexting, and the creation of deepfakes. The author reflects on the double-edged impact of advancements like Microsoft’s VALL-E and image synthesis models like DALL-E 2, drawing a trajectory of the evolving threat landscape in social engineering through deepfakes and exploiting human cognitive biases.
@@ 3.3 Automated hacking
PentestGPT [41] or GPTs [42], which are the custom versions of ChatGPT that can be created for a specific purpose like GP(en)T(ester) [43]. Pentest Reporter [44] are introduced as applications built on ChatGPT, designed to assist in penetration testing, which is a sanctioned simulation of cyberattacks on systems to evaluate security. However, these tools could also be adapted for malicious purposes in automated hacking. Many emerging tools, such as WolfGPT, XXXGPT, and WormGPT, have been invented; however, no study has yet evaluated and compared their real offensive capabilities. Gupta et al. [32] noted that an AI model could scan new code for similar weaknesses with a comprehensive dataset of known software vulnerabilities, pinpointing potential attack vectors. While AI-assisted tools like PentestGPT are intended for legitimate and constructive uses, there is potential for misuse by malicious actors who could create similar models to automate unethical hacking procedures.
If fine-tuned to identify vulnerabilities, craft exploitation strategies, and execute those strategies, these models could potentially pose significant threats to cybersecurity. However, this enormous task should be divided into smaller segments, such as reconnaissance, privilege escalation, and more. Temara [45] outlines how ChatGPT can be utilized during the reconnaissance phase by employing a case study methodology to demonstrate collecting reconnaissance data such as IP addresses, domain names, network topologies, and other critical information like SSL/TLS cyphers, ports and services, and operating systems used by the target. Happe et al. [46] investigate the use of LLMs in Linux privilege escalation. The authors introduce a benchmark for automated testing of LLMs’ abilities to perform privilege escalation using a variety of prompts and strategies. They implement a tool named Wintermute, a Python program that supervises and controls the privilege-escalation attempts to evaluate different models and prompt strategies. Their findings indicate that GPT-4 generates the highest quality commands and responses. In contrast, Llama2-based models struggle with command parameters and system descriptions. In some scenarios, GPT-4 achieved a 100% success rate in exploitation.
@@ 3.4 Attack payload generation
Some studies [32, 47] have highlighted the capacity of LLMs, particularly ChatGPT, for payload generation. Our examination of GPT-4’s current abilities confirmed its proficiency in generating payloads and embedding them into PDFs (as an example) using a reverse proxy (Fig. 10). The following is a summation of the frameworks GPT-4 utilizes with successful payload code generation, accompanied by their respective primary functions:
@@ 3.5 Malware Code Generation
Gupta et al. [32] mentioned they could obtain potential ransomware code examples by utilizing a DAN jailbreak. We tested all the existing DAN techniques outlined in [48]. At the time of our research, these techniques were no longer functional; therefore, we could not reproduce samples of WannaCry, Ryuk, REvil, or Locky, as addressed by [32]. However, we generated an educational ransomware code, as shown in Fig. 11, applying basic code obfuscation like renaming and control flow flattening. ChatGPT has garnered significant attention from the cybersecurity community, leading to the implementation of robust filters. Nonetheless, this does not imply that other models, such as the Chinese 01.ai[49], will have an equivalent opportunity to mitigate the potential for misuse in generating malicious code.
@@ 3.6 Polymorphic malware
The usage of LLMs could see the rise of malware, which integrates improved evasion techniques and polymorphic capabilities [50]. This often relates to overcoming both signature detection and behavioural analysis. An LLM-based malware agent could thus focus on rewriting malware code, which could change the encryption mode used or produce obfuscated code which is randomized for each build [51]. Gupta et al. [32] outlined a method of getting ChatGPT to seek out target files for encryption and thus mimic ransomware behaviour, but where it mutated the code to avoid detection. They even managed to embed a Python interpreter in the malware where it could query ChatGPT for new software modules.
@@ 4 Cyber Defence
In the ever-evolving cybersecurity battlefield, the “Cyber Defence” segment highlights the indispensable role of GenAI in fortifying digital fortresses against increasingly sophisticated cyber threats. This section is dedicated to exploring how GenAI technologies, renowned for their advanced capabilities in data analysis and pattern recognition, are revolutionizing the approaches to cyber defence. Iturbe et al. [54] described the AI4CYBER framework. This framework includes AI4TRIAGE (methods to perform alert triage to determine the root cause of an attack), AI4VUN (identifies vulnerabilities), AI4FIX ( test for vulnerabilities and automatically fix them), and I4COLLAB (privacy-aware information-sharing mechanism).
@@ 4.2 Cybersecurity Reporting
Using LLMs provides a method of producing Cyber Threat Intelligence (CTI) using NLP techniques. For this, Perrina et al. [57] created the Automatic Generation of Intelligence Reports (AGIR) system to link text data from many data sources. For this, they found that AGIR has a high recall value (0.99) without any hallucinations, along with a high score of the Syntactic Log-Odds Ratio (SLOR).
@@ 4.3 Threat Intelligence
Bayer et al. [58] address the challenge of information overload in the gathering of CTI from open-source intelligence. A novel system is introduced, utilizing transfer learning, data augmentation, and few-shot learning to train specialized classifiers for emerging cybersecurity incidents. In parallel, Microsoft Security Copilot [59] has been providing CTI to its customers using GPT, and operational use cases have been observed, such as the Cyber Dome initiative in Israel [60].
@@ 4.4 Secure Code Generation and Detection
Machine learning in code analysis for cybersecurity has been elaborated very well [61]. Recent progress in NLP has given rise to potent language models like the GPT series, encompassing LLM like ChatGPT and GPT-4 [62]. Traditionally, SAST is a method that employs Static Code Analysis (SCA) to detect possible security vulnerabilities. We are interested in seeing whether SAST or GPT could be more efficient in decreasing the vulnerability window. The window of vulnerability is defined as when the most vulnerable systems apply the patch minus the time an exploit becomes active. The precondition is met if two milestones that assume the detection of vulnerabilities verify their effectiveness, along with the vendor patch [63].
Laws in some countries, like China, ban the reporting on zero-days (see articles 4 and 9 of [64]), and contests like the Tianfu Cup [65], which is a systematic effort to find zero days, proliferate zero-day discovery continuously. Therefore, this precondition may not be satisfied timely, especially if the confirmation of vulnerabilities is not verified. A wide window of vulnerability threatens national security if a zero-day has been taken against critical infrastructures. DARPA introduces an important challenge that may help overcome the threat (AIxCC) [66]). Moreover, this topic touches a part of the BSI studies [67, 68], and where we can define two main classifications of software testing for cybersecurity bugs:
  • SAST: This is often called White Box Testing, which is a set of algorithms and techniques used for analyzing source code. It operates automatically in a non-runtime environment to detect vulnerabilities such as hidden errors or poor source code during development.
  • Dynamic Application Security Testing (DAST): This follows the opposite approach and analyzes the program while it operates. Functions are called with values in the variables as each line of code is checked, and possible branching scenarios are guessed. Currently, GPT-4 and other LLMs cannot provide DAST capabilities because the code needs to run within the runtime for this to work, requiring many deployment considerations.
@@ 4.5 Vulnerability Detection and Repair
Using GitHub data for code synthesis, Mark et al. [73] presented Codex, which is a significant advancement in GPT language models. GitHub Copilot functions on the basis of this improved model. When assessed on the HumanEval dataset, designed to gauge the functional accuracy of generating programs based on docstrings, Codex achieved a remarkable f28.8% success rate. GPT-J obtained an 11.4% success rate, whereas GPT-3 produced a 0% success rate. One notable finding was that the model performed better with repeated sampling; given 100 samples per problem, the success rate increased to 70.2%. Even with these encouraging outcomes, Codex still has certain drawbacks. It particularly struggles with complex docstrings and variable binding procedures. The article discusses the wider consequences of using such powerful code-generation technologies, including safety, security, and financial effects.
Cheshkov et al. [74] discovered in a technical assessment that the ChatGPT and GPT-3 models, although successful in other code-based tasks, were only able to match the performance of a dummy classifier for this specific challenge. Utilizing a dataset of Java files sourced from GitHub repositories, the study emphasized the models’ current limitations in the domain of vulnerability detection. However, the authors remain optimistic about the potential of future advancements, suggesting that models like GPT-4, with targeted research, could eventually make significant contributions to the field of vulnerability detection.
@@
4.6 Evaluating GenAI for Code Security
Ten security issues were introduced by the OWASP top 10 for LLMs [76]. They are as follows: Prompt Injection, Unauthorized Output Processing, Training Data Poisoning, Denial of Service Attacks, Supply Chain Security Flaws, Disclosure of Sensitive Information, Unauthorized Plugin Development, Abnormal Agency, Overdependence, and Model Theft.
Elgedawy et al. [77] analysed the ability of LLM to produce both secure and insecure code and conducted experiments using GPT-3.5, GPT-4, Google Bard and Google Gemini from Google. This involved nine basic tasks in generating code and assessing for functionality, security, performance, complexity, and reliability. They found that Bard was less likely to link to external libraries and, thus, was less exposed to software chain issues. There were also variable levels of security and code integrity, such as input validation, sanitization, and secret key management, and while useful for automated code reviews, LLMs often require manual reviews, especially in understanding the context of the deployed code. For security, GPT-3.5 seemed to be more robust for error handling and secure coding practices when there is security consciousness applied to the prompt, there was a lesser focus on this with GPT-4, but where there were more advisory notes given. Overall, Gemini produced the most code vulnerabilities, and the paper advised users to be careful when deploying secure code from Gemini.
@@ 4.7 Developing Ethical Guidelines
Kumar et al. [78] outlined the ethical challenges related to LLMs and where the datasets that they were trained on could be open to breaches of confidentiality, including five major threats: prompt injection, jailbreaking, personally identifiable information (PII) exposure, sexually explicit content, and hate-based content. They propose a model that provides an ethical framework for scrutinizing the ethical dimensions of an LLM within the testing phase. The MetaAID framework [79] focuses on strengthening cybersecurity using Metaverse cybersecurity Q&A and attack simulation scenarios, along with addressing concerns around the ethical implications of user input. The framework is defined across five dimensions:
  • Ethics: This defines an alignment with accepted moral and ethical principles.
  • Legal Compliance: Any user input does not violate laws and/or regulations. This might relate to privacy laws and copyright protection.
  • Transparency: User inputs must be clear in requirements and do not intend to mislead the LLM.
  • Intent Analysis: User input should not have other intents, such as jailbreaking the LLM.
  • Malicious intentions: User input should be free of malicious intent, such as performing hate crimes.
  • Social Impact: This defines how user input could have a negative effect on society, such as searching for ways to do harm to others, such as related to crashing the stock market or planning a terrorist attack.
@@ 4.8 Incident Response and Digital Forensics
Using a pre-trained LLM for artefact comprehension, evidence searching, code development, anomaly identification, incident response, and education was examined by Scanlon et al. [24]. Expert specialization is still needed for many other applications, including low-risk ones. The main areas of strength include assurance, inventiveness, and avoiding the blank page problem, particularly in areas where ChatGPT excels, including creating forensic scenarios and providing evidence assurance. However, caution must be used to prevent ChatGPT hallucinations. Code generation and explanations, including creating instructions for tool integration that can serve as a springboard for further research, are another useful application.
Regarding weaknesses, Scanlon discovered that having a high-quality, current training model was crucial; in the absence of one, ChatGPT’s analysis could be prejudiced and out of date. In general, if it is trained on comparatively old data, it may not be able to locate the most recent artefacts. Furthermore, ChatGPT’s accuracy decreases with job specificity, and its accuracy is further diminished when analyzing non-textural input, including network packets. Another issue with some evidence logs was their length, which frequently required prefiltering before analysis. The output of ChatGPT is frequently not predictable, which makes it inappropriate for reproducibility, which is the last issue found.